# What’s the Weight?

### v1

@@ 4.4 Summary Takeaways
In summary, the simulation settings described above are designed to vary the associations between $Y$, $A$, $X$, and $S$ with increasing complexity, in order to introduce either/both confounding and/or selection bias into the data generating mechanism. In settings with neither confounding nor selection bias, all existing and proposed methods perform comparably well, estimating the ACD with low bias and MSE, and nominal coverage. In settings of confounding bias (without selection bias), the methods which account for both $A$ and $X$, namely all of the multiple regression and IPTW-based estimators, tend to perform comparatively well, whereas methods that only model $A$ suffer from high bias and negligible coverage. When $X$ is introduced into the generation of either/both $S$ and $A$, that is, when selection or selection + confounding bias exists, we see that standard analytic approaches lead to biased estimates and do not attain nominal coverage, whereas the proposed methods perform empirically well, supporting theoretical results derived in the previous section. Lastly, our sensitivity analysis reveals that the proposed methods which rely on contrasts and inverse probability weighting, rather than outcome modeling and direct standardization, are more robust to misspecification, particularly as the degree of effect heterogeneity between the groups of interest increases.
@@ 5 NHANES: Race, SES, and Telomere Length
We now turn our attention to our motivating example, which comes from the National Health and Nutrition Examination Survey (NHANES). NHANES is a nationally representative survey sample collected by the Centers for Disease Control, which contains a wealth of demographic, dietary, physical examination, laboratory, and questionnaire data on individuals. Our variable of interest is the survey participant’s self-reported race. Racial groups are categories of people treated as a social entity by virtue of physical characteristics. While an individual’s race is often self-reported, racial groups may represent commonalities such as geographic origin, genetic ancestry, cultural norms and traditions, social history, or current social circumstances (Williams et al., 2010). From a statistical perspective, race is often a proxy for other sources of shared variation which must be accounted for when making inferential statements. We focus on non-Hispanic participants who self-identify as Black or White as, particularly in the US, it is known that Black individuals tend to have poorer health outcomes than White individuals due to structural inequalities (Arias and Xu, 2022).
@@ 5.1 Motivating Question
Telomeres are regions of DNA at the ends of chromosomes that protect against cellular aging and senescence (Lin et al., 2016). Telomere shortening is associated with disease incidence and a variety of health outcomes (Shammas, 2011). Variation in telomere lengths is attributable to age and shortening during mitosis, in addition to genetics, social history, exposure to environmental contaminants, and psycho-social stressors (Jiang et al., 2007; Blackburn et al., 2015). Paradoxically, Black individuals typically have longer telomeres than White individuals (Needham et al., 2013; Venkateswaran et al., 2024), with plausible explanations citing leukocyte cell composition, population stratification, or genetic differences (Freedman et al., 1997; Lin et al., 2016; Hansen et al., 2016). Others posit that environmental exposures associated with longer telomere length underlie this so-called ‘race effect,’ as these exposures differ significantly between Blacks and Whites (Needham et al., 2020; Roberts et al., 2022). Further, studies have found telomere length to be comparable between Black and White individuals in socio-economically homogeneous populations (Geronimus et al., 2015), calling into question purely genetic explanations. Our motivating question is this: if we could hypothetically balance socioeconomic status (SES; $X$) between Black and White individuals ($A$) in a nationally representative sample ($S$), would we still see significant Black/White differences in telomere length ($Y$)?
Figure 2: Simulation setup and results. Columns denote each of eight simulation settings. The top row gives the assumed relationships between variables for each setting. The second row gives the parameter settings for each data generating mechanism. The third row gives the average percent bias and Monte Carlo standard errors for each method across 200 simulation replicates. The fourth row gives the coverage probabilities for each method.
@@ 5.2 NHANES Sampling Design
The stratified, clustered sampling design of the NHANES survey is designed to over-represent Black individuals and underrepresented groups in particular, to achieve a more nationally representative sample. Namely, NHANES samples in four stages. In the first stage, primary sampling units (PSUs; i.e., counties) are selected. Counties are stratified and selected with probability proportional to their weighted average population counts, where the weights give a higher probability of selection to counties with a higher proportion of underrepresented subgroups (e.g., certain age groups, marginalized racial groups, lower-income households). In the second stage, the PSUs are further divided (e.g., into city blocks) and sub-sampled, and in the third stage, households are sampled from these sample segments. Again, households within segments that have a greater proportion of underrepresented age, ethnic, or income groups are given additional weight for being selected. Lastly, individuals within households are sampled according to designated age-sex-race/ethnicity screening subdomains. This complex sampling scheme motivates our proposed methods, as sample selection depends on self-reported race and potential confounders (i.e., socioeconomic status).
@@ 5.3 Study Sample
The 1999-2002 NHANES waves consisted of 7,839 participants with recorded laboratory measures, representing our survey sample $(S=1)$. Our primary endpoint, $Y$, was the log-transformed mean ratio of an individual’s telomere length to a standard reference DNA sample across all leukocyte cell types (mean T/S). Telomere length was assayed via quantitative polymerase chain reaction (PCR; Cawthon, 2002; Needham et al., 2013). We focus on the 1999-2002 NHANES waves, as they featured 4-year adjusted survey weights designed for aggregating data across cohorts. Among the initial 7,839 participants, 5,308 (67.7%) self-identified as either non-Hispanic White or non-Hispanic Black, the levels of our group variable $A$. Excluding those participants without our outcome of interest, our final analytic sample contained 5,298 Non-Hispanic White or Non-Hispanic Black identifying participants with measured telomere length. As discussed previously, race was our variable of interest. We further adjusted our models for study participant age, sex, and blood cell composition, including complete blood count and five part differential consisting of the proportion of neutrophils, eosinophils, basophils, lymphocytes, and monocytes, to account for known differences in these factors (Freedman et al., 1997), as well as twelve indicators of SES, collectively representing our confounding variables $X$. Ten of these, namely marital status, education level, household income, insurance status, Special Supplemental Nutrition Program for Women, Infants, and Children (WIC) usage, household size, home ownership, home type, food security status, and an individual’s poverty income ratio (PIR), were extracted directly from the NHANES demographic and occupation questionnaires. Occupation category was constructed by mapping occupation group codes in the NHANES 1999-2002 occupation questionnaire to the national statistics socioeconomic job classifications, and employment status was derived from three occupational measures: type of work done last week, hours worked last week at all jobs, and main reason for not working last week (Rehkopf et al., 2008; Rose et al., 2005). In later modeling, we perform complete case analysis on $n=5,270$ participants after removing 28 individuals without measured blood composition.
@@
Descriptive statistics for our analytic sample are presented in Table LABEL:tab:nhanesdesc, both overall and stratified by self-reported race. As shown, non-Hispanic Black participants had a median telomere length (mean T/S ratio) of 1.05 (i.e., 0.05 times longer than the standard reference DNA), as compared to non-Hispanic White participants with a median telomere length of 0.98. With respect to the identified socioeconomic characteristics, non-Hispanic Black versus White participants differed significantly across all 12 SES indicators, with Black participants having lower educational attainment, lower household income, and lower home ownership rates, as well as higher food insecurity and WIC utilization rates, larger household sizes, and a higher proportion uninsured, unemployed, and below the federal poverty line. Further, a higher proportion of Black participants reported not working or working in low blue collar occupations, a higher proportion reported being widowed, divorced, or separated, and a higher proportion reported residing in apartments, all of which suggest lower socio-economic status (Rose et al., 2005). Based on these factors, we then calculated propensity scores with race as the outcome using unweighted and survey-weighted logistic regressions. The results from these models fit are given in supplementary Table LABEL:tab:nhanespropscore. The propensity scores from these models were then carried forward to our later outcome models in order to estimate the ACD of telomere length by race.
@@ 5.4 Results
We compared our proposed approaches to six methods that one might consider in practice, each of which successively increase in their respective adjustments for potential confounding and selection bias. Namely, we fit a multiple linear regression (Approach 2 in Table 1) of log-transformed telomere length on race, age, sex, the blood composition variables, and the socioeconomic indicators, not accounting for the potential confounding or selection bias in the study. We then fit a survey-weighted multiple regression ((Approach 4), and inverse probability of treatment weighted (IPTW) multiple regression (Approach 5), a regression weighted by the product of the within sample propensity score weights and the survey weights (Approach 6), and a regression weighted by the product of the survey-weighted propensity score weights and the survey weights (Approach 7), not unlike our simulation study. We omitted the oracle regression, simple linear regression, and the naive g-computation methods, as they were not applicable (as with the oracle regression) or inadvisable (as with the simple linear regression and naive g-computation). For the proposed methods, standard errors for the ACD were calculated via Equation (9), whereas for the existing methods, we calculate either model-based or design-based standard errors where appropriate.
@@
ACD estimates and corresponding 95% confidence intervals across the various analytic approaches are presented in Table 3 below. In general, the ACD estimates across different analytic strategies tend to attenuate in magnitude as we made more appropriate adjustments for confounding and selection bias. For example, the multiple linear regression had an ACD estimate of 0.0265, corresponding to approximately a 2.7% longer telomere length, on average, among non-Hispanic Black individuals as compared to non-Hispanic White individuals, adjusting for all other factors (95% CI: 0.0106-0.0423). In contrast the ACD point estimates from our proposed approaches varied from 0.0132 (IPW2) to 0.0176 (OM), though these average controlled differences were not statistically different from zero. These results suggest that there is a strong confounding relationship between SES and race. Moreover, the methods which properly incorporated the NHANES stratified, clustered design tended to have more conservative standard error estimates than those which did not, as evidenced by the wider confidence intervals (Table 3).

### v2

@@ 4.4 Summary Takeaways
In summary, the simulation and sensitivity analyses described above vary the associations between $Y$, AA, $\boldsymbol{X}$, $\boldsymbol{U}$, and $S$ to introduce varying degrees of confounding and/or selection bias into the data generating mechanism. In our main simulation, across all eight simulation scenarios, the drivers of performance are (1) whether the estimator properly adjusts for the source(s) of bias (confounding, selection, or both), and (2) whether its variance calculation accounts for the stratified, clustered survey design and any weight variability. In settings without confounding or selection bias, most existing and proposed methods perform similarly well, providing estimates of the ACD with low bias, low MSE, and nominal coverage. In settings with confounding bias (but no selection bias), methods that account for both $A$ and $\boldsymbol{X}$, such as multiple regression and IPTW-based estimators, tend to perform well. Methods that only model $A$ often show high bias and negligible coverage. When $\boldsymbol{X}$ is introduced into the generation of either/both $S$ and $A$, that is, when selection or selection and confounding bias exists, we see that standard analytic approaches lead to biased estimates and do not attain nominal coverage, whereas the proposed methods perform empirically well, supporting theoretical results derived in the previous section.
Our sensitivity analyses reveal several key points that have implications in practice. First, in settings of good propensity score overlap, all proposed methods are unbiased with nominal coverage, outcome modeling is most efficient, IPW2 most variable, and AIPW intermediate. In poor overlap scenarios, outcome modeling and AIPW remain relatively stable, while IPW1 and IPW2 have larger bias or variance. Researchers should examine the degree of overlap and possible violations of the positivity assumption when conducting analyses. Second, when either the outcome or propensity model alone is mildly or moderately misspecified, AIPW and IPW1/IPW2 maintain low bias and nominal coverage. In contrast, outcome modeling incurs bias and low coverage under outcome model misspecification, and IPW2 can be unstable under propensity model misspecification. When both models are misspecified, AIPW is the most robust. Third, when the selection mechanism is unknown, estimating $\Pr(S=1\mid A,\boldsymbol{X})$ may lead to slight increases in bias and under coverage for IPW1/IPW2, but AIPW remains robust. Lastly, even when key sampling covariates are omitted from the estimated selection model, IPW1, IPW2, and AIPW maintains lower bias and good coverage, while outcome modeling shows modest negative bias but retains high coverage due to conservative SEs.
Thus, in real-world survey analyses, where both confounding and selection mechanisms may be complex or partially unknown, overlap can be limited, and model misspecification is a concern, the doubly robust augmented IPW estimator provides the best balance of bias protection, valid variance estimation under complex sampling, and efficiency. We therefore recommend our proposed AIPW estimator as the default analytic approach for estimating controlled differences or causal contrasts in complex survey settings such as the one described in our motivating example.
@@ 5 NHANES: Race, SES, and Telomere Length
We now present our motivating example, which comes from NHANES. NHANES is a nationally representative survey sample collected by the Centers for Disease Control, which contains a wealth of demographic, dietary, physical examination, laboratory, and questionnaire data on individuals. Our group variable of interest is the survey participant’s self-reported race. Racial groups are categories of people treated as a social entity by virtue of physical characteristics. While an individual’s race is often self-reported, racial groups may represent commonalities such as geographic origin, genetic ancestry, cultural norms and traditions, social history, or current social circumstances (Williams et al., 2010). From a statistical perspective, race is often a proxy for other sources of shared variation which must be accounted for when making inferential statements. We focus on non-Hispanic participants who self-identify as Black or White as, particularly in the US, it is known that Black individuals tend to have poorer health outcomes than White individuals due to structural inequalities (Arias and Xu, 2022).
@@ 5.1 Motivating Question
Telomeres are regions of DNA at the ends of chromosomes that protect against cellular aging and senescence (Lin et al., 2016). Telomere shortening is associated with disease incidence and many health outcomes (Shammas, 2011). Variation in telomere lengths is attributable to age and shortening during mitosis, in addition to genetics, social history, exposure to environmental contaminants, and psychosocial stressors (Jiang et al., 2007; Blackburn et al., 2015). Paradoxically, Black individuals typically have longer telomeres than White individuals (Needham et al., 2013; Venkateswaran et al., 2024), with plausible explanations citing leukocyte cell composition, population stratification, or genetic differences (Freedman et al., 1997; Lin et al., 2016; Hansen et al., 2016). Others posit that environmental exposures associated with longer telomere length underlie this so-called ‘race effect,’ as these exposures differ significantly between Blacks and Whites (Needham et al., 2020; Roberts et al., 2022). Further, studies have found telomere length to be comparable between Black and White individuals in socioeconomically homogeneous populations (Geronimus et al., 2015), calling into question purely genetic explanations. Our motivating question is this: if we could hypothetically balance SES ($\boldsymbol{X}$) between Black and White individuals ($A$) in a nationally representative sample ($S$), would we still see significant Black/White differences in telomere length ($Y$)?
@@ 5.2 NHANES Sampling Design
The stratified, clustered sampling design of the NHANES survey is designed to over-represent Black individuals and underrepresented groups in particular, to achieve a more nationally representative sample. Namely, NHANES samples in four stages. In the first stage, primary sampling units (PSUs; i.e., counties) are selected. Counties are stratified and selected with probability proportional to their weighted average population counts, where the weights give a higher probability of selection to counties with a higher proportion of underrepresented subgroups (e.g., certain age groups, marginalized racial groups, lower-income households). In the second stage, the PSUs are further divided (e.g., into city blocks) and subsampled, and in the third stage, households are sampled from these sample segments. Again, households within segments that have a greater proportion of underrepresented age, ethnic, or income groups are given additional weight for being selected. Lastly, individuals within households are sampled according to designated age-sex-race/ethnicity screening subdomains. This complex sampling scheme motivates our proposed methods, as sample selection depends on self-reported race and potential confounders (i.e., SES).
@@ 5.3 Study Sample
The 1999-2002 NHANES waves consisted of 7,839 participants with recorded laboratory measures, representing our survey sample $(S=1)$. Our primary endpoint, $Y$, was the log-transformed mean ratio of an individual’s telomere length to a standard reference DNA sample across all leukocyte cell types (mean T/S). Telomere length was assayed via quantitative polymerase chain reaction (PCR; Cawthon, 2002; Needham et al., 2013). We focus on the 1999-2002 NHANES waves, as they featured 4-year adjusted survey weights designed for aggregating data across cohorts. Among the initial 7,839 participants, 5,308 (67.7%) self-identified as either non-Hispanic White or non-Hispanic Black, the levels of our group variable $A$. Excluding those participants without our outcome of interest, our final analytic sample contained 5,298 Non-Hispanic White or Non-Hispanic Black identifying participants with measured telomere length. As discussed previously, race was our variable of interest. We further adjusted our models for study participant age, sex, and blood cell composition, including complete blood count and five part differential consisting of the proportion of neutrophils, eosinophils, basophils, lymphocytes, and monocytes, to account for known differences in these factors (Freedman et al., 1997), as well as twelve indicators of SES, collectively representing our confounding variables $\boldsymbol{X}$. Ten of these, namely marital status, education level, household income, insurance status, Special Supplemental Nutrition Program for Women, Infants, and Children (WIC) usage, household size, home ownership, home type, food security status, and an individual’s poverty income ratio (PIR), were extracted directly from the NHANES demographic and occupation questionnaires. Occupation category was constructed by mapping occupation group codes in the NHANES occupation questionnaire to the national statistics socioeconomic job classifications, and employment status was derived from three occupational measures: type of work done last week, hours worked last week at all jobs, and main reason for not working last week (Rehkopf et al., 2008; Rose et al., 2005). In later modeling, we perform complete case analysis on $n=5,270$ participants after removing 28 individuals without measured blood composition.
@@
Full descriptive statistics for our analytic sample are presented in Supplemental Table C1, both overall and stratified by self-reported race. Non-Hispanic Black participants had a median telomere length (mean T/S ratio) of 1.05 (i.e., 0.05 times longer than the standard reference DNA), as compared to non-Hispanic White participants with a median telomere length of 0.98. With respect to the identified socioeconomic characteristics, non-Hispanic Black versus White participants differed significantly across all 12 SES indicators, with Black participants having lower educational attainment, lower household income, and lower home ownership rates, as well as higher food insecurity and WIC utilization rates, larger household sizes, and a higher proportion uninsured, unemployed, and below the federal poverty line. Further, a higher proportion of Black participants reported not working or working in low blue collar occupations, a higher proportion reported being widowed, divorced, or separated, and a higher proportion reported residing in apartments, all of which suggest lower socioeconomic status (Rose et al., 2005). Based on these factors, we then calculated propensity scores with race as the outcome using unweighted and survey-weighted logistic regressions. The results from these models fit are given in Supplemental Table C2, and plots of the resulting covariate balance statistics are given in Supplemental Figures C1 and C2. As shown, the SES covariate balance was improved across all variables after adjustment by both the unweighted and survey-weighted propensity scores, bringing all below a threshold of 0.1 for both the absolute standardized mean differences and Kolmogorov-Smirnov statistics. The propensity scores from these models were then carried forward to our later outcome models in order to estimate the ACD of telomere length by race.
@@ 5.4 Results
We compared our proposed approaches to six methods that one might consider in practice, each of which successively increase in their respective adjustments for potential confounding and selection bias. Namely, we fit a multiple linear regression (Approach 2 in Table 1) of log-transformed telomere length on race, age, sex, the blood composition variables, and the socioeconomic indicators. We then fit a survey-weighted multiple regression (Approach 4), and inverse probability of treatment weighted (IPTW) multiple regression (Approach 5), a regression weighted by the product of the within sample propensity score weights and the survey weights (Approach 6), and a regression weighted by the product of the survey-weighted propensity score weights and the survey weights (Approach 7), not unlike our simulation study. We omitted the oracle regression and simple linear regression, as they were not applicable (as with the oracle regression) or inadvisable (as with the simple linear regression). We calculated standard errors for the ACD with the proposed methods via Equation (9), whereas for the existing methods, we calculate model- versus design-based standard errors where appropriate.
@@
ACD estimates and corresponding 95% confidence intervals across the various analytic approaches are presented in Table 3 below. In general, the ACD estimates across different analytic strategies tend to attenuate in magnitude as we made more appropriate adjustments for confounding and selection bias. For example, the multiple linear regression had an ACD estimate of 0.0265, corresponding to approximately a 2.7% longer telomere length, on average, among non-Hispanic Black individuals as compared to non-Hispanic White individuals, adjusting for all other factors (95% CI: 0.0106-0.0423). In contrast, the ACD point estimates from our proposed approaches varied from 0.0122 (AIPW) to 0.0176 (OM), though these ACD estimates were not statistically different from zero. These results suggest that there is a strong confounding relationship between SES and race. Moreover, the methods which incorporated the NHANES stratified, clustered design tended to have more conservative standard error estimates, as evidenced by the wider confidence intervals (Table 3).
Method ACD Estimate 95% Confidence Interval
Multiple Regression 0.0265 0.0106, 0.0424
IPTW Estimator 0.0262 0.0079, 0.0445
Survey-Weighted Multiple Regression 0.0298 -0.0010, 0.0606
IPTW Multiple Regression 0.0183 0.0053, 0.0312
IPTW + Survey-Weighted Multiple Regression 0.0220 -0.0088, 0.0527
Weighted IPTW + Survey-Weighted Multiple Regression 0.0186 -0.0126, 0.0498
Proposed Outcome Modeling and Direct Standardization 0.0176 -0.0029, 0.0381
Proposed Inverse Probability Weighting 1 0.0141 -0.0134, 0.0416
Proposed Inverse Probability Weighting 2 0.0131 -0.0081, 0.0342
Proposed Augmented Inverse Probability Weighting 0.0122 -0.0077, 0.0321